Plot a graph of sinx


1
Expert's answer
2021-11-03T08:05:46-0400
"f(x)=\sin x"

Domain: "(-\infin, \infin)"

Range: "[-1, 1]"

There are no asymptotes.


"f(-x)=\sin(-x)=-\sin x=-f(x), x\in \R"

The function is odd. The graph is symmetric with respect to the origin.


"f(x+T)=f(x+2\pi)=f(x), x\in \R"

The function is periodic. The period is "T=2\pi."


"y" ​- intersection:"x=0, f(0)=\sin(0)=0."

Point "(0, 0)."

The graph passes through the origin.

"x" ​- intersection(s): "y=0, 0=\sin x=>x=\pi n, n\in \Z."

Points "(\pi n, 0), n\in \Z"



"f'(x)=(\sin x)=\cos x"

Find the critical number(s)


"f'(x)=0=>\cos x=0=>x=\dfrac{\pi}{2}+\pi k, k\in \Z"

If "-\dfrac{\pi}{2}+2\pi k<x<\dfrac{\pi}{2}+2\pi k, k\in \Z," then "f'(x)>0, f(x)" increases.


If "\dfrac{\pi}{2}+2\pi k<x<\dfrac{3\pi}{2}+2\pi k, k\in \Z," then "f'(x)<0, f(x)" decreases.


"f(-\dfrac{\pi}{2}+2\pi k)=\sin(-\dfrac{\pi}{2}+2\pi k)=-1,k\in \Z"


"f(\dfrac{\pi}{2}+2\pi k)=\sin(\dfrac{\pi}{2}+2\pi k)=1,k\in \Z"


The function "f(x)=\sin x" has the local maximum with value of "1" at "x=\dfrac{\pi}{2}+2\pi k, k\in \Z."


The function "f(x)=\sin x" has the local minimum with value of "-1" at "x=-\dfrac{\pi}{2}+2\pi k, k\in \Z."



"f''(x)=(\cos x)'=-\sin x"

Find the point(s) of inflection


"f''(x)=0=>-\sin x=0=>x=\pi m, m\in \Z""f(\pi m)=\sin(\pi m)=0,m\in \Z"

Points of inflection: "(\pi m, 0), m\in\Z"

If "-\pi+2\pi m<x<2\pi m, m\in \Z," then "f''(x)>0, f(x)" is concave up.


If "2\pi m<x<\pi+2\pi m, m\in \Z," then "f''(x)<0, f(x)" is concave down.


Sketch the graph of "f(x)=\sin x"